# thanks: corresponding author

### v3

@@ 4 Projected Gradient Descent with Guaranteed Convergence
([88, Lemma 26]) Let $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ be an MPO with MPO ranks $(r_{1},\dots,r_{n-1})$. Denote by $\underline{\sigma}({\bm{\rho}^{\star}})$ the smallest TT singular value of $\bm{\rho}^{\star}$ [^15] Specifically, we reshape $\bm{\rho}$ into an $n$-th order tensor $\mathcal{X}$ of size $d^{2}\times d^{2}\times\cdots\times d^{2}$ and define the $\ell$-th unfolding matrix of tensor $\mathcal{X}$ as $\bm{X}^{\langle\ell\rangle}\in\mathbb{C}^{d^{2\ell}\times d^{2n-2\ell}}$, where the $(s_{1}\cdots s_{\ell},s_{\ell+1}\cdots s_{n})$-th element of $\bm{X}^{\langle\ell\rangle}$ is given by $\bm{X}^{\langle\ell\rangle}(s_{1}\cdots s_{\ell},s_{\ell+1}\cdots s_{n})=\mathcal{X}(s_{1},\dots,s_{n})$. With the $\ell$-th unfolding matrix $\mathcal{X}^{\langle\ell\rangle}$ and the MPO ranks, we can obtain its smallest TT singular value $\underline{\sigma}(\mathcal{X})=\min_{\ell=1}^{n-1}\sigma_{r_{\ell}}(\mathcal{X}^{\langle\ell\rangle})$.. Then for any perturbation $\bm{E}\in\mathbb{C}^{d^{n}\times d^{n}}$ with $C_{n}||\bm{E}||_{F}\leq\underline{\sigma}({\bm{\rho}^{\star}})$ for some constant $C_{n}\geq 500n$, we have
@@
$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}<\frac{\underline{\sigma}({\bm{\rho}^{\star}})(d-1)(1-\delta)^{2}}{600n(1+\delta^{2}+(4d-2)\delta)},$ (24)
@@
$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg{(}\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$ (25)
@@
The proof is given in Appendix E. Theorem 4 provides a linear convergence analysis of the PGD algorithm along with its error characterization. Moreover, the recovery error presented in (27) aligns closely with the results in Theorems 1 and 2. Since the problem (22) is nonconvex (as the MPO set is nonconvex), Theorem 4 requires an initialization close to the target state that satisfies (24). [^16] The right-hand side of (24) depends on $\underline{\sigma}({\bm{\rho}^{\star}})$, which could potentially be exponentially small for high-rank states. However, this is not the case for many practical states, such as those at low temperatures or systems close to pure states. However, as will be illustrated in Section 5, our experiments show that even a randomly initialized PGD can efficiently converge to a stable recovery error. To relax the requirement on the initialization, the following result establishes another convergence guarantee with an additional requirement on the projection.
@@
$\|\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F}^{2}\leq(1+b)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2},$
$\text{with}\ b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}.$ (26)
@@
$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg{(}\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$ (27)
@@
[…]
$\cdot O\bigg{(}\frac{d\overline{r}\sqrt{(1+\delta)n\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}+\delta\|\bm{\rho}^{\star}\|_{F}\bigg{)}.$
Its proof is given in Appendix F. The scaling term $\sqrt{n-1}$ arises due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)=\mathcal{P}_{\operatorname{trace}=1}(\text{SVD}_{\bm{r}}^{tt}(\cdot))$. Recall that $\gamma_{t}(\bm{\rho}^{\star})=1$ for all $\bm{\rho}^{\star}$ when spherical 3-design POVMs (i.e., $t\geq 3$), and $\gamma_{t}(\bm{\rho}^{\star})=\gamma(\bm{\rho}^{\star})$ for spherical 2-design POVMs (i.e., $t=2$). Theorem 5 shows that when $\gamma_{t}(\bm{\rho}^{\star})$ is small—true for all MPOs when using $3$-designs and for those having relatively uniformed probability distribution with 2-design POVMs—the spectral initialization provides a relatively good estimation with a polynomially large $M$. In particular, considering exact $t$-design POVMs or cases where $\delta$ is small, (29) ensures that $\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}$ scales in ther order of $\frac{n^{2}d^{2}\overline{r}^{2}}{M}$ when ignoring the log terms and $\gamma_{t}(\bm{\rho}^{\star})$. Note that the requirements on the number of measurements are slightly more stringent than those in Theorem 1 and Theorem 2 due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)$. However, as will be illustrated in Section 5, our numerical experiments show that even a randomly initialized PGD can very efficiently find an estimate of the target state. A theoretical explanation for this phenomenon will be the focus of future research.
@@
5 Numerical Experiments with Local IC-POVMs
@@
$\bigg{\{}\begin{bmatrix}\frac{1}{2}&0\\ 0&0\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}\\ \frac{\sqrt{2}}{6}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{2\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{2\pi}{3}}&\frac{1}{6}\end{bmatrix},$
$\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{4\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{4\pi}{3}}&\frac{1}{6}\end{bmatrix}\bigg{\}}.$ (30)
@@
While our main results (Theorems 1 and 2) do not apply to such a measurement setting, we show that certain MPO states can be recovered using this IC-POVM using a number of state copies only linear in $n$ and the PGD and PSGD algorithms developed in Section 3. A more systematic study on what types of MPO states can be recovered efficiently using the local IC-POVM can be found in [98], using a different optimization algorithm based on maximal likelihood estimation (MLE). As an example, here we consider a random matrix product density operator (MPDO) state [68] of the following form (1), i.e, the $(i_{1}\cdots i_{n},j_{1}\cdots j_{n})$-th entry of $\bm{\rho}^{\star}$ can be generated by
@@
where the symbol * represents the complex conjugate, $\otimes$ denotes the Kronecker product, $K_{\ell}$ is an arbitrary positive integer that controls purity of the generated state, and each matrix ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ has dimension $\kappa\times\kappa$, except for ${\bm{A}_{1}^{\star}}^{i_{1},a_{1}}$ and ${\bm{A}_{n}^{\star}}^{i_{n},a_{n}}$, which are of dimension $1\times\kappa$ and $\kappa\times 1$ respectively. Each entry of ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ is i.i.d. randomly generated with both its real and imaginary parts drawn uniformly from the interval $[-1,1]$. In the experiments, we set $K_{\ell}=10$ to ensure the state is sufficiently mixed; using $K_{\ell}=1$ would generate a pure state. The generated $\bm{\rho}^{\star}$ is an MPO with a bond dimension of $\overline{r}=\kappa^{2}$ and is always PSD. To be a physical state, $\bm{\rho}^{\star}$ also needs to have unit trace, which can be done by calculating
@@
[…]
6 Conclusion
@@ Appendix A Proof of Lemma 1
$=$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\bigg{(}\bigg{(}\sum_{i=1}^{d^{2n}}\langle\bm{A}_{i},\bm{\rho}\rangle\bigg{)}^{2}-\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\bigg{)}$
@@ Appendix B Proof of LABEL:{l2_norm_of_approximate_2-designs_RIP}
[…]
@@ Appendix C Proof of Theorem 1
$\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle=\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle\leq\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle,$ (37)
@@
$\begin{split}\overline{\mathbb{X}}_{\bm{r}}=\Big{\{}&\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\ \bm{\rho}=\bm{\rho}^{\dagger},\|\bm{\rho}\|_{F}\leq 1,\operatorname{trace}(\bm{\rho})=0,\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},\\ &\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}},\ell\in[n-1],r_{0}=r_{n}=1\Big{\}}.\end{split}$ (38)
@@
The rest of the proof is to bound $\max_{\bm{\rho}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle$, which will be achieved by using a covering argument. First, when conditioned on $\bm{A}_{k}$, we consider any fixed value of $\bm{\rho}^{(p)}=[\bm{X}_{1}^{(p_{1})},\dots,\bm{X}_{n}^{(p_{n})}]\in\overline{\mathbb{X}}_{2\bm{r}}$ and according to [33, Appendix B-A], construct a $\xi$-net $\{\bm{\rho}^{(1)},\ldots,\bm{\rho}^{(N_{1}\cdots N_{n})}\}=\{[\bm{X}_{1}^{(1)},\dots,\bm{X}_{n}^{(1)}],\dots,[\bm{X}_{1}^{(N_{1})},\dots,\bm{X}_{n}^{(N_{n})}]\}$ such that $\sup_{\bm{X}_{\ell}:\|\bm{X}_{\ell}\|_{F}\leq 1}\leavevmode\nobreak\ \min_{p_{\ell}\leq N_{\ell}}\|\bm{X}_{\ell}-\bm{X}_{\ell}^{(p_{\ell})}\|_{F}\leq\xi,\ell=1,\dots,n$ with covering number $\Pi_{\ell=1}^{n}N_{\ell}\leq(\frac{4+\xi}{\xi})^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}$.

### v4

@@ IV Projected Gradient Descent with Guaranteed Convergence
[…]
@@
$\hskip-28.45274pt\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}<\frac{\underline{\sigma}({\bm{\rho}^{\star}})(d-1)(1-\delta)^{2}}{600n(1+\delta^{2}+(4d-2)\delta)},$ (24)
@@
$\hskip-22.76228pt\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg(\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg)$ (25)
@@
The proof is given in Appendix E. Theorem 4 provides a linear convergence analysis of the PGD algorithm along with its error characterization. Moreover, the recovery error presented in (27) aligns closely with the results in Theorems 1 and 2. Since the problem (22) is nonconvex (as the MPO set is nonconvex), Theorem 4 requires an initialization close to the target state that satisfies (24). [^16] The right-hand side of (24) depends on $\underline{\sigma}({\bm{\rho}^{\star}})$, which could potentially be exponentially small for high-rank states. However, this is not the case for many practical states, such as those at low temperatures or systems close to pure states. However, as will be illustrated in Section V, our experiments show that even a randomly initialized PGD can efficiently converge to a stable recovery error. To relax the requirement on the initialization, the following result establishes another convergence guarantee with an additional requirement on the projection.
@@
$\|\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F}^{2}\leq(1+b)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2},\text{with}\ b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}.$ (26)
@@
$\hskip-22.76228pt\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg(\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg)$ (27)
@@
$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}\leq(1+\sqrt{n-1})\cdot O\bigg(\frac{d\overline{r}\sqrt{(1+\delta)n\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}+\delta\|\bm{\rho}^{\star}\|_{F}\bigg).$ (29)
Its proof is given in Appendix F. The scaling term $\sqrt{n-1}$ arises due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)=\mathcal{P}_{\operatorname{trace}=1}(\text{SVD}_{\bm{r}}^{tt}(\cdot))$. Recall that $\gamma_{t}(\bm{\rho}^{\star})=1$ for all $\bm{\rho}^{\star}$ when spherical 3-design POVMs (i.e., $t\geq 3$), and $\gamma_{t}(\bm{\rho}^{\star})=\gamma(\bm{\rho}^{\star})$ for spherical 2-design POVMs (i.e., $t=2$). Theorem 5 shows that when $\gamma_{t}(\bm{\rho}^{\star})$ is small—true for all MPOs when using $3$-designs and for those having relatively uniformed probability distribution with 2-design POVMs—the spectral initialization provides a relatively good estimation with a polynomially large $M$. In particular, considering exact $t$-design POVMs or cases where $\delta$ is small, (29) ensures that $\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}$ scales in ther order of $\frac{n^{2}d^{2}\overline{r}^{2}}{M}$ when ignoring the log terms and $\gamma_{t}(\bm{\rho}^{\star})$. Note that the requirements on the number of measurements are slightly more stringent than those in Theorem 1 and Theorem 2 due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)$. However, as will be illustrated in Section V, our numerical experiments show that even a randomly initialized PGD can very efficiently find an estimate of the target state. A theoretical explanation for this phenomenon will be the focus of future research.
@@
V Numerical Experiments with Local IC-POVMs
@@
$\bigg\{\begin{bmatrix}\frac{1}{2}&0\\ 0&0\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}\\ \frac{\sqrt{2}}{6}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{2\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{2\pi}{3}}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{4\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{4\pi}{3}}&\frac{1}{6}\end{bmatrix}\bigg\}$. (30)
@@
While our main results (Theorems 1 and 2) do not apply to such a measurement setting, we show that certain MPO states can be recovered using this IC-POVM using a number of state copies only linear in $n$ and the PGD and PSGD algorithms developed in Section III. A more systematic study on what types of MPO states can be recovered efficiently using the local IC-POVM can be found in [98], using a different optimization algorithm based on maximal likelihood estimation (MLE). As an example, here we consider a random matrix product density operator (MPDO) state [68] of the following form (1), i.e, the $(i_{1}\cdots i_{n},j_{1}\cdots j_{n})$-th entry of $\bm{\rho}^{\star}$ can be generated by
@@
where the symbol * represents the complex conjugate, $\otimes$ denotes the Kronecker product, $K_{\ell}$ is an arbitrary positive integer that controls purity of the generated state, and each matrix ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ has dimension $\kappa\times\kappa$, except for ${\bm{A}_{1}^{\star}}^{i_{1},a_{1}}$ and ${\bm{A}_{n}^{\star}}^{i_{n},a_{n}}$, which are of dimension $1\times\kappa$ and $\kappa\times 1$ respectively. Each entry of ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ is i.i.d. randomly generated with both its real and imaginary parts drawn uniformly from the interval [1,1][-1,1]. In the experiments, we set $K_{\ell}=10$ to ensure the state is sufficiently mixed; using $K_{\ell}=1$ would generate a pure state. The generated $\bm{\rho}^{\star}$ is an MPO with a bond dimension of $\overline{r}=\kappa^{2}$ and is always PSD. To be a physical state, $\bm{\rho}^{\star}$ also needs to have unit trace, which can be done by calculating
@@
$\text{PGD:}\ O\bigg(\underbrace{4^{n}n{\overline{r}}^{3}}_{\text{inner product}}+\underbrace{4^{n}{\overline{r}}^{2}}_{\text{TT-SVD}}+\underbrace{n{\overline{r}}^{2}}_{\text{unit trace}}\bigg),$
$\text{PSGD:}\ O\bigg(\frac{N}{B}(\underbrace{Bn{\overline{r}}^{3}}_{\text{inner product}}+\underbrace{n({\overline{r}}+B)^{3}}_{\text{rounding}}+\underbrace{n{\overline{r}}^{2}}_{\text{unit trace}})\bigg).$
VI Conclusion
@@ Appendix A Proof of Lemma 1
$\!\!\!\!=\!\!\!\!$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\big(\big(\sum_{i=1}^{d^{2n}}\langle\bm{A}_{i},\bm{\rho}\rangle\big)^{2}-\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\big)$
@@ Appendix B Proof of LABEL:{l2_norm_of_approximate_2-designs_RIP}
$\bigg\langle\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w},\frac{d^{2n}}{K}\bm{\rho}^{\otimes 2}\bigg\rangle$ $\!\!\!\!=\!\!\!\!$ $\frac{d^{n}(\operatorname{trace}(\bm{\rho}^{\otimes 2})+\operatorname{trace}(\bm{S}\bm{\rho}^{\otimes 2}))}{K(d^{n}+1)}$ (36)
@@ Appendix C Proof of Theorem 1
$\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ $\!\!\!\!=\!\!\!\!$ $\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ (37)
$\!\!\!\!\leq\!\!\!\!$ $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle,$
@@
$\begin{split}\overline{\mathbb{X}}_{\bm{r}}=\Big\{&\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\ \bm{\rho}=\bm{\rho}^{\dagger},\|\bm{\rho}\|_{F}\leq 1,\operatorname{trace}(\bm{\rho})=0,\\ &\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},\\ &\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}},\ell\in[n-1],r_{0}=r_{n}=1\Big\}.\end{split}$ (38)
@@
The rest of the proof is to bound $\max_{\bm{\rho}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle$, which will be achieved by using a covering argument. First, when conditioned on $\bm{A}_{k}$, we consider any fixed value of $\bm{\rho}^{(p)}=[\bm{X}_{1}^{(p_{1})},\dots,\bm{X}_{n}^{(p_{n})}]\in\overline{\mathbb{X}}_{2\bm{r}}$ and according to [33, Appendix B-A], construct a $\xi$-net $\{\bm{\rho}^{(1)},\ldots,\bm{\rho}^{(N_{1}\cdots N_{n})}\}=\{[\bm{X}_{1}^{(1)},\dots,\bm{X}_{n}^{(1)}],\dots,[\bm{X}_{1}^{(N_{1})},\dots,\bm{X}_{n}^{(N_{n})}]\}$ such that $\sup_{\bm{X}_{\ell}:\|\bm{X}_{\ell}\|_{F}\leq 1}~\min_{p_{\ell}\leq N_{\ell}}\|\bm{X}_{\ell}-\bm{X}_{\ell}^{(p_{\ell})}\|_{F}\leq\xi,\ell=1,\dots,n$ with covering number $\Pi_{\ell=1}^{n}N_{\ell}\leq(\frac{4+\xi}{\xi})^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}$.